If f(x)=x+4 and g(x)=2x^2-x-1, find the composition (fog)(2) please I could use some help.
thank you

Answer by stanbon(75887)

If f(x)=x+4 and g(x)=2x^2-x-1, find the composition (fog)(2)
---------------
(fog)(2) = f[g(2)] = f[2*2^2-2-1] = f[5]= 5+4 = 9
========================
Cheers,
stan H.